OpenStudy (anonymous):

HELP ASAP. MEDAL GIVEN Solve the equation below for x: log(5x) + log(2x) =1 x = 10/7 x = 1 x = 1; x = -1 There is no solution.

OpenStudy (solomonzelman):

At first lets use the identity, \[\color{blue}{Log(a)+Log(b)=Log(a \times b) }\] so the left side equals \[\log(5x \times 2x)\]so, so far we have\[\log(10x^2)=1\] can you solve it from here?

OpenStudy (anonymous):

Honestly No... -_-

OpenStudy (solomonzelman):

\[\color{red}{ \log(10x^2)=1 }\]\[\color{red}{ 2\log(10x)=1 }\]\[\color{red}{ \log(10x)=1/2 }\]what about now?

OpenStudy (anonymous):

Umm, I honestly don't know what any of that is suppose to mean.. im not that bright in math

OpenStudy (solomonzelman):

Oh nvm I mad a mistake

OpenStudy (solomonzelman):

\[\color{orangered}{\log(10)+\log(x^2)=1 }\] when base is not specified, it's equal to 10. \[\color{orangered}{\log_{10}(10)+\log_{10}(x^2)=1 }\]\[\color{orangered}{1+\log_{10}(x^2)=1 }\]\[\color{orangered}{\log_{10}(x^2)=0 }\] you should be able to do it from here.

OpenStudy (anonymous):

1 ?

OpenStudy (solomonzelman):

\[\log_{10}(x^2)=0\]\[2\log_{10}(x)=0\]\[\log_{10}(x)=0\]\[x=1\]

OpenStudy (solomonzelman):

you are right!

OpenStudy (anonymous):

thanks :)

OpenStudy (solomonzelman):

You Welcome!